please make it clear how i open the brackets best thanx
17-(13x-(15-19x)-21x)=25-((11x-23)-(18x-7))

Answer by sabanasir(22)

17-(13x-(15-19x)-21x)=25-((11x-23)-(18x-7))
=> 17 - (13x - 15 + 19x - 21x) = 25 - (11x - 23 - 18x + 7)
=> 17 - 13x + 15 - 19x + 21x = 25 - 11x + 23 + 18x - 7
=> 17 + 15 - 13x - 19x + 21x = 25 + 23 - 7 - 11x + 18x
=> 32 -32x + 21x = 48 - 7 + 7x
=> 32 - 11x = 41 + 7x
=> 32 - 41 = 11x + 7x
=> - 9 = 18x
=> - 9 / 18 = x
=> - 1/2 = x